# Searching for new physics with 136Xe double beta decay spectrum in PandaX-4T

### v1

@@
Searching for new physics with 136Xe double beta decay spectrum in PandaX-4T
@@
Xiangdong JiDepartment of Physics, University of Maryland, College Park, Maryland 20742, USA
@@ Abstract
We present a comprehensive analysis of the 136Xe spectrum utilizing kgyr of 136Xe exposure from the PandaX-4T experiment.
The analysis yields the most precise measurement to date of the 136Xe two-neutrino double beta decay () half-life, years, the uncertainty of which is reduced by a factor of two compared to our previous result.
@@
Double beta decay () is a rare nuclear process [DBD:origin].
@@
In contrast, if no neutrinos are emitted, the process is referred to as neutrinoless double beta decay (), which can only occur if neutrinos are Majorana particles [NLDBD:probability, NLDBD:review1, NLDBD:review2].
Additionally, if one or two bosons are emitted in processes, they are typically referred to as “Majorons” [Majoron:goldstone, Majoron:Left-handed, Majoron:Unconventional].
Although has been experimentally observed and measured in several studies, such as those involving 136Xe [EXO-200:Xe136dbd2014, KamLAND-Zen:kesi31, NEXT:Xe136dbd, PandaX:2022kwg], neither nor Majoron-emitting has been detected.
Recent experiments have established the most stringent lower limits on the half-life of different isotopes, such as 136Xe [KamLAND:Xenldbd2023, EXO-200:2019nldbd], 76Ge [GERDA:Genldbd2020, Majorana:2023Genldbd, CDEX:2024Genldbd], and 130Te [CUORE:2022Tenldbd].
The half-life is linked to the effective Majorana neutrino mass by [NLDBD:review2]
@@
The current understanding of NME remains incomplete, with significant discrepancies observed in different nuclear many-body approaches and potential influences of quenching effects [NME:Status, Xe136nldbd:2015review].
Since and share the same initial and final nuclear states, accurately determining NME is essential to constrain NME predictions [kesi31:theory].
A refined expression for the half-life is presented as follows [kesi31:theory]:
where is the effective axial-vector coupling constant.
In contrast to the standard phase space factor , the phase space factor accounts for the dependence on lepton energies from the energy denominators of NMEs.
The parameter is defined as , where the subleading NME is only sensitive to contributions from the lightest states of the intermediate nucleus, while the NME is also sensitive to higher-lying states.
The experimentally accessible parameter , which can be measured through electron energy spectrum fits, allows for the discrimination between different theoretical models and thus provides new insights into NME calculations.
Majorons are potential dark matter candidates and play roles in various cosmological and astrophysical processes [Majoron:darkmatter1, Majoron:darkmatter2, Majoron:darkmatter3].
@@
Different modes may produce identical summed electron energy spectra, which are classified into four types based on the spectral index 1, 2, 3 and 7 [Majoron:n137, Majoron:n2].
Previous experimental searches have established stringent constraints on these processes with 136Xe [KamLAND-Zen:Majoron, EXO-200:Majoron2021].
Previous searches for and Majoron-emitting of 136Xe have been limited to relatively high energy thresholds, specifically between 500 keV and 800 keV [KamLAND-Zen:kesi31, KamLAND-Zen:Majoron, EXO-200:Majoron2021].
@@
The PandaX-4T detector, which uses a natural xenon target and was initially designed for dark matter searches in the low-energy region, has expanded its detection capability to the MeV range [PandaX:2022kwg, PandaX-4T:Xe136nldbd].
In this paper, we leverage the nearly complete spectrum to perform the measurement of and the searches for Majoron-emitting of 136Xe, using a dataset of the commissioning run (Run0) and the first science run (Run1) of the PandaX-4T experiment [PandaX:2025WIMP, PandaX:SuperWIMP, PandaX-4T:Xe136nldbd] with a total 136Xe exposure of kgyr.
The energy region of interest (ROI) for the spectrum fit is from 20 keV to 2800 keV.
The PandaX-4T detector is a dual-phase time projection chamber (TPC) containing 3.7 tonnes of natural xenon in its active volume.
@@
More detailed description of the detector can be found in Ref. [PandaX-4T:2021bab].
@@
Energy reconstruction utilizes the area of and .
Spatial reconstruction determines the vertical (z axis) position via drift velocity and the time delay between the and signals, while the horizontal (x-y axis) coordinates are obtained through maximum likelihood estimation of the charge distribution pattern of observed in the top PMT array.
The data production and event selection procedures follow those described in Refs. [PandaX-4T:Xe136nldbd, PandaX:SuperWIMP, PandaX:2024med, PandaX:2023ggs].
The fiducial volume (FV) from the previous study [PandaX:SuperWIMP], defined as the innermost cleanest part of the detector, is adopted in this analysis.
@@
The uncertainty arises from the LXe density and the difference between the geometrically-calculated volume and the volume scaled by the [PandaX:SuperWIMP].
The potential shift between the reconstructed energy and the true energy as well as the energy resolution are modelled using the 164 keV peak (from Xe), along with the 236 keV peaks (from Xe and Xe), the 1460 keV peak (from 40K), and the 2615 keV peak (from 232Th), obtained from data in the control region outside of the FV.
This ensures that the energy response model is completely uncorrelated with the data used in the final fit.
Five parameters are used to model the energy response, independently for Run0 and Run1.
The energy resolution is modeled as a Gaussian function, with the width given by , where is the reconstructed energy in keV.
The residual energy nonlinearity is modeled as , where is the true energy.
The calibrated values and uncertainties (Table 1), along with the covariance matrix , are used to fit the Run0 and Run1 data.
The specific approach is consistent with that in Ref. [PandaX:SuperWIMP].
@@
denotes the 5-parameter detector response model (see text), with its means and uncertainties determined from mono-energetic peaks obtained from the peripheral data of FV.
The total detector efficiency comprises three components: the single-site (SS) fraction, quality cut (QC) efficiency, and ROI acceptance, estimated with methods inherited from Ref. [PandaX:SuperWIMP] and summarized in Table 1.
The SS fraction is defined as the ratio of SS to the sum of SS and multi-site (MS) energy deposition events in the detector.
We validate the SS fraction within the ROI using 232Th calibration data together with 232Th events simulated with the BambooMC simulation framework [Chen:2021asx].
In this analysis, the QC efficiency of Run0 and Run1 are ()% and ()% respectively.
To generate the signal spectra for the fit, the theoretical spectra of and in Eq. 2 are calculated from the numerical wave functions which are the solutions of Dirac equation [NLDBDtheory:leftright], obtained with the package RADIAL [Diracwave] by assuming a uniform distribution of the nuclear charge inside the nucleus.
@@
For and Majoron-emitting spectra, we use the Decay0 package [BxDecay0] to generate events with momentum of two electrons, and then simulate the detector response with BambooMC package as well.
@@
Background events originate from three primary sources: liquid xenon, external structure, and solar neutrinos, as summarized in Table 2.
124Xe and 125I are taken from the measurements reported in Ref. [PandaX:2024xpq].
For xenon isotopes in Run0, including 127Xe, 129mXe and 131mXe, we adopt the xenon-evolution analysis of Ref. [PandaX:SuperWIMP].
All remaining xenon isotopes are left free in the fit.
85Kr concentration is determined via - cascade tagging through the isomeric 85mRb [Collon:2004xs].
@@
The external structure, including both the detector materials and the stainless steel platform (SSP), contributes to the background.
Activities of 232Th, 238U, 60Co, and 40K in the detector materials are taken from measurements using high-purity germanium (HPGe) counting stations [PandaX:materialHGe], while activities of 232Th and 238U in SSP are taken from Ref. [PandaX:SSP].
The solar and 7Be neutrinos contributions are taken from Refs. [Chen:2016eab, BOREXINO:2014pcl].
@@
The fitted counts are obtained from the 136Xe half-life measurement, the measurement and the Majoron-emitting search with mode .
@@
Assuming no new physics is observed in this study, a one-dimensional binned likelihood fit to measure the half-life of 136Xe is performed first, as shown in Fig. 2.
@@
The Gaussian penalty term of the energy response contains the five-parameter and the covariant matrix in Run-.
The Gaussian penalty terms are used to constrain the nuisance parameters , , and , which represent the relative uncertainties in the overall efficiency, the signal selection, and background model (see Table 2), respectively.
The radioactivities of external structures, 124Xe concentrations, and solar and 7Be fluxes are identical in Run0 and Run1.
@@
Figure 2: The SS data spectra and the fit for the 136Xe half-life are shown for Run0 (top) and Run1 (bottom) from 20 keV to 2800 keV with a bin size of 4 keV. The horizontal axis represents the reconstructed energy in the data. Xe∗ includes the contributions from 124Xe, 125Xe, 127Xe, 129mXe, 131mXe, and 133Xe. The lower panel shows the residuals together with , and bands.
@@
The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), as well as the bias of the signal selection is (Run0) and (Run1).
@@
It is worth noting that this result is slightly smaller than our previous result from Run0 data [PandaX:2022kwg], due to the use of the updated and slightly smaller 136Xe abundance of [PandaX:SSP], compared to the previous value of .
@@
Two-dimensional profile likelihood scan for 136Xe rate and .
The black dot indicates the best-fit value of .
The contours correspond to the 68.3, 90, 95.4, and 99.7 CL.
The likelihood is almost the same as Eq. 3, but with a different definition of in Run- which is modified as
where corresponds to the PDF corresponding to the first term of the Taylor expansion of the decay rate of , and corresponds to the PDF corresponding to the second term.
The parameter , which comprises , and in Eq. 2 as well as the total detection efficiencies, varies linearly with .
The rest of the parameters remain the same as in Eq. 4.
The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), as well as the bias of the signal selection is (Run0) and (Run1).
We obtained a best fit of , consistent with theoretical predictions [kesi31:theory] and the KamLAND-Zen results [KamLAND-Zen:kesi31].
Fig. 3 shows the two-dimensional confidence intervals for the 136Xe rate and , revealing only a slight positive correlation.
The inclusion of the second-order contribution is found to have a negligible impact on the 136Xe decay rate.
The 136Xe decay half-life is estimated to be years.
This result is consistent with the 136Xe -only fit, when the second-order contribution was not considered.
Figure 4: Spectra fits with Majoron-emitting decay signals ().
The figure is the same as Fig. 2 but with the upper limits (90% CL) of signals illustrated as hatched histograms. The lower panel shows the residuals together with , and bands.
In the fit of Majoron-emitting , the likelihood is similar to Eq. 3, except that 136Xe is moved from the signal term to the background term, and then Majoron-emitting is introduced as the signal.
@@
The extension of the lower bound of the ROI from 600 keV (as in EXO-200 [EXO-200:Majoron2021]) to 20 keV considerably improves the sensitivity to mode , where the Majoron-emitting spectrum peaks at approximately 600 keV, leading to the most stringent constraint.
However, our sensitivity to modes 1, 2, and 3 is limited due to our 136Xe exposure being only 0.16 times that of EXO-200 [EXO-200:Majoron2021] and our higher background level.
@@
The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), as well as the bias of the signal selection is (Run0) and (Run1).
Table 3: The 90% CL lower limits on half-lives for different Majoron-emitting decay models and the comparison to EXO-200 results [EXO-200:Majoron2021]
@@
Our measurement of is consistent with theoretical predictions and the results from KamLAND-Zen.
@@
This project is supported in part by grants from National Key R&D Program of China (Nos. 2023YFA1606200, 2023YFA1606202), National Science Foundation of China (Nos. 12090060, 12090062, 12305121, U23B2070), and by Office of Science and Technology, Shanghai Municipal Government (grant Nos. 21TQ1400218, 22JC1410100, 23JC1410200, ZJ2023-ZD-003). We thank for the support by the Fundamental Research Funds for the Central Universities. We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Thomas and Linda Lau Family Foundation, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.

### v2

# Precise 136Xe Double Beta Decay Measurement in PandaX-4T with Implications on the Nuclear Matrix Elements and Majorons

@@
Precise 136Xe Double Beta Decay Measurement in PandaX-4T with Implications on the Nuclear Matrix Elements and Majorons
@@
Xiangdong JiState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Department of Physics, University of Maryland, College Park, Maryland 20742, USA
@@ Abstract
We present a comprehensive analysis of the 136Xe spectrum utilizing of 136Xe exposure from the PandaX-4T experiment.
The analysis yields the most precise measurement to date of the 136Xe two-neutrino double beta decay () half-life, years, the uncertainty of which is reduced by a factor of 2 compared to our previous result.
@@
Double beta decay () is a rare nuclear process [25].
@@
In contrast, if no neutrinos are emitted, the process is referred to as neutrinoless double beta decay (), which can occur only if neutrinos are Majorana particles [20, 39, 5, 26].
Additionally, if one or two bosons are emitted in processes, they are typically referred to as “Majorons” [16, 23, 24].
Although has been experimentally observed and measured in several studies, such as those involving 136Xe [7, 22, 35, 40], neither nor Majoron-emitting has been detected.
Recent experiments have established the most stringent lower limits on the half-life of different isotopes, such as 136Xe [1, 8], 76Ge [4, 9, 44, 2], and 130Te [3].
The half-life is linked to the effective Majorana neutrino mass by [26]
@@
The current understanding of NME remains incomplete, with significant discrepancies observed in different nuclear many-body approaches and potential influences of quenching effects [19, 18].
Since and share the same initial and final nuclear states, accurately determining NME is essential to constrain NME predictions [41].
A refined expression for the half-life, retaining terms up to second order in the Taylor expansion, is presented as follows [41]:
where is the effective axial-vector coupling constant and denotes the leading-order Gamow-Teller component of the NME.
In contrast to the standard phase space factor , the phase space factor accounts for the dependence on lepton energies of the energy denominators of NMEs.
The parameter is defined as , where the subleading NME is sensitive only to contributions from the lightest states of the intermediate nucleus, while the NME is also sensitive to higher-lying states.
The experimentally accessible parameter , which can be measured through electron energy spectrum fits, allows for the discrimination between different theoretical models, and thus, provides new insights into NME calculations.
Specifically, a zero (nonzero) value of implies the higher (single) state dominance hypothesis of [41].
Majorons are potential dark matter candidates and play roles in various cosmological and astrophysical processes [28, 10, 29].
@@
Different modes may produce identical summed electron energy spectra, which are classified into four types based on the spectral index 1, 2, 3, and 7 [27, 34].
Previous experimental searches have established stringent constraints on these processes with 136Xe [21, 6].
Previous measurements of and searches for Majoron-emitting of 136Xe have been limited to relatively high energy thresholds, specifically between 500 keV and 800 keV [22, 21, 6].
@@
The PandaX-4T detector, which uses a natural xenon target and was initially designed for dark matter searches in the low-energy region, has expanded its detection capability to the MeV range [40, 45].
In this Letter, we leverage the nearly complete spectrum to perform the measurement of and the searches for Majoron-emitting modes of of 136Xe, using a dataset of the commissioning run (Run0, 94.8 days of live time) and the first science run (Run1, 163.5 days) of the PandaX-4T experiment [12, 30, 45] with a total 136Xe exposure of .
The energy region of interest (ROI) for the spectrum fit is from 20 keV to
2800 keV.
The PandaX-4T detector is a dual-phase time projection chamber (TPC) containing 3.7 tons of natural xenon in its active volume.
@@
A more detailed description of the detector can be found in Ref. [33].
@@
Energy reconstruction utilizes the charge of and , where stands for detected by the bottom PMT array, following the procedure detailed in Ref. [45, 30, 32, 43].
Spatial reconstruction determines the vertical (-axis) position via drift velocity and the time delay between the and signals, while the horizontal (- and -axes) coordinates are obtained through maximum likelihood estimation of the charge distribution pattern of observed in the top PMT array [30].
The data production and event selection procedures follow those described in Refs. [45, 30, 32, 43].
The fiducial volume (FV) from the previous study [30], defined as the innermost, cleanest part of the detector, is adopted in this analysis.
@@
The uncertainty arises from the LXe density and the difference between the geometrically calculated volume and the volume scaled by the uniformly distributed calibration source [30].
Compared with our previous study of the 136Xe decay [40], the present analysis extends the analysis window down to 20 keV.
Similar to Ref. [30], monoenergetic peaks at 41.5 keV (from ), 164 keV (Xe), and 236 keV (127Xe and Xe) are used for energy calibration, together with high-energy peaks from 137Cs, 60Co, 40K, and the 232Th decay chain.
The energy response is modeled using five parameters, which account for a possible shift between the reconstructed and true energies as well as the energy resolution [30].
The energy resolution is parametrized by a Gaussian function with a width given by
, where is the reconstructed energy in keV. The residual energy nonlinearity is modeled as , where denotes the true energy.
The calibrated parameter values and their uncertainties, (Table 1), together with the corresponding covariance matrix , are used to fit the Run0 and Run1 data independently.
Due to the lack of sufficient high-energy calibration data within the fiducial volume, the energy response model is derived from physics data in a control region outside the FV, providing a closely matched but statistically uncorrelated dataset for the final fit [45].
@@
denotes the 5-parameter detector response model (see text), with its means and uncertainties determined from monoenergetic peaks obtained from the control data outside of the FV.
Sources
Run0 (94.8 d)
Run1 (163.5 d)
Detectorresponse
[keV-1]
[keV]
Overallefficiency
SS fraction ()
Quality cut
Signalselection
LXe density [g/cm3]
136Xe abundance
FV mass [kg]
The total detection efficiency comprises three components: the single-site (SS) fraction, quality cut (QC) efficiency, and ROI acceptance, estimated with methods inherited from Ref. [30] and summarized in Table 1.
The SS fraction is defined as the ratio of SS to the sum of SS and multisite (MS) energy deposition events in the detector.
We validate the SS fraction within the ROI using 232Th calibration data together with 232Th events simulated with the BambooMC simulation framework [15].
In this analysis, the QC efficiency of Run0 and Run1 are ()% and ()%, respectively.
To generate the signal spectra for the fit, the theoretical spectra of the first and second terms of the Taylor expansion in Eq. (2) are calculated from the numerical wave functions, which are the solutions of the Dirac equation [42], obtained with the package RADIAL [38] by assuming a uniform distribution of the nuclear charge inside the nucleus.
@@
For and Majoron-emitting spectra, we use the Decay0 package [36] to generate events with momenta of two electrons, and then, simulate the detector response with the BambooMC package as well.
@@
Background events originate from three primary sources: liquid xenon, the detector material and the stainless steel platform (SSP) material, and solar neutrinos, as summarized in Table 2.
124Xe and 125I are taken from the measurements reported in Ref. [13].
For 127Xe, 129mXe, and 131mXe in Run0, we adopt the xenon-evolution analysis from Ref. [30].
In Run1, due to the absence of neutron calibration, the activities of these xenon isotopes are very low, and their time evolution cannot be determined precisely.
Therefore, we allow their activities to float in the fit.
The activities of 125Xe and 133Xe are left free in Run0 and are negligible in Run1.
85Kr concentration is determined via - cascade tagging through the isomeric 85mRb [17].
@@
The material of the detector and the SSP contributes to the background.
Activities of 232Th, 238U, 60Co, and 40K in the detector components are taken from measurements using high-purity germanium counting stations [37], while activities of 232Th and 238U in SSP components are taken from Ref. [31].
The solar and 7Be neutrino contributions are taken from Refs. [14, 11].
@@
The fitted counts are obtained from the 136Xe half-life measurement, the measurement, and the Majoron-emitting search with mode .
@@
Assuming no new physics is observed in this study, a one-dimensional binned likelihood fit is first performed to measure the half-life of 136Xe , using a Frequentist approach, as shown in Fig. 2.
@@
The Gaussian penalty term of the energy response contains the five-parameter and the covariance matrix in Run-.
The Gaussian penalty terms are used to constrain the nuisance parameters , , and , which represent the relative deviations in the overall efficiency, the signal selection, and the background model (see Table 2), respectively.
The radioactivities of detector and SSP materials, 124Xe concentrations, and solar and 7Be neutrino fluxes are identical in Run0 and Run1.
@@
Figure 2: The SS data spectra and the fit for the 136Xe half-life are shown for Run0 (top) and Run1 (bottom) from 20 keV to 2800 keV with a bin size of 4 keV. The horizontal axis represents the reconstructed energy in the data. Xe∗ denotes the contributions from 124Xe, 125Xe, 127Xe, 129mXe, 131mXe, and 133Xe. Material includes the contributions from SSP 232Th, SSP 238U, detector 60Co, detector 40K, detector 232Th, and detector 238U. The lower panel shows the residuals together with , , and bands.
@@
The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), and the bias of the signal selection is (Run0) and (Run1).
@@
It is worth noting that this result is slightly smaller than our previous result from Run0 data [40], due to the use of the updated and slightly smaller 136Xe abundance of [31], compared to the previous value of .
@@
The likelihood is almost the same as Eq. (3), but with a different definition of in Run-, which is modified as
where and denote the normalized energy spectrum values in the bin for the first and second terms of the Taylor expansion of the decay rate, respectively.
The parameter is defined as , where and denote the total detection efficiencies for the theoretical spectra of the first and second terms of the Taylor expansion, respectively.
The rest of the parameters remain the same as in Eq. (4).
The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), and the bias of the signal selection is (Run0) and (Run1).
We obtained a best fit of .
Figure 3 compares the measured values of and the 136Xe half-life obtained by KamLAND-Zen [22] and by this work.
Although both results are consistent with , our measured value has the opposite sign to that reported by KamLAND-Zen ().
Assuming fully uncorrelated uncertainties, the two measurements differ by approximately .
Our result is consistent with quasiparticle random-phase approximation predictions [22] at the level and remains compatible with shell-model calculations [22] within .
However, the current precision does not allow us to exclude conclusively and prevents us from discriminating between the higher-state dominance and single-state dominance hypotheses [41].
Achieving such discrimination will require improved precision in future experiments.
@@
Comparison of 136Xe half-lives and values from the KamLAND-Zen experiment [22] and our present analysis.
The gray dashed line indicates the zero .
The inclusion of the second-order contribution is found to have a negligible impact on the extracted 136Xe decay rate.
The corresponding half-life is determined to be years, which is consistent with the result obtained from a 136Xe -only fit that neglects the second-order contribution.
Figure 4: Spectra fits with Majoron-emitting decay signals () for Run0 (top) and Run1 (bottom), respectively.
The figures are similar to Fig. 2 but with the upper limit (90% CL) on the Majoron-emitting signal (mode ) illustrated as hatched histograms. The lower panel shows the residuals together with , , and bands.
In the fit of Majoron-emitting , the likelihood is similar to Eq. (3), except that 136Xe is moved from the signal term to the background term, and then, Majoron-emitting is introduced as the signal.
@@
The extension of the lower bound of the ROI from 600 keV (as in EXO-200 [6]) to 20 keV considerably improves the sensitivity to mode , where the Majoron-emitting spectrum peaks at approximately 600 keV, leading to the most stringent constraint.
However, our sensitivity to modes 1, 2, and 3 is limited due to our 136Xe exposure being only 0.16 times that of EXO-200 [6] and our higher background level.
@@
The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), and the bias of the signal selection is (Run0) and (Run1).
Table 3: The 90% CL lower limits on half-lives for different Majoron-emitting decay models and the comparison with EXO-200 results [6]
@@
Our measurement of is consistent with theoretical predictions.
The values from KamLAND-Zen and this work are broadly consistent, with a modest (1.7) difference.
@@
Acknowledgements—This project is supported in part by grants from the National Key R&D Program of China (No. 2023YFA1606200 and No. 2023YFA1606202), National Science Foundation of China (No. 12090060, No. 12090062, No. 12305121, No. U23B2070), and by Office of Science and Technology, Shanghai Municipal Government (Grants No. 21TQ1400218, No. 22JC1410100, No. 23JC1410200, No. ZJ2023-ZD-003). We are grateful for the support of the Fundamental Research Funds for the Central Universities. We also thank for their sponsorship: the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Thomas and Linda Lau Family Foundation, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.
Data availability—The data that support the findings of this article are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.